# Interpreting the Win Ratio in Hierarchical Composite Endpoints: Challenges, Limitations, and Perspectives with Examples from Chronic Kidney Disease Trials

### v1

@@ 3 Variance dependence
In this section we consider the impact of the variance of the continuous random variables on the win statistics when one or more continuous components are present in the HCE, both from an estimand and estimation perspectives.
@@ 3.1 The impact of the overall variance
Let $Y_{0}\sim\mathcal{N}(\mu_{0},\sigma_{0}^{2})$ and $Y_{1}\sim\mathcal{N}(\mu_{1},\sigma_{1}^{2})$ be two independent normally distributed random variables, where $Y_{i}$ denotes the response under treatment $i=0,1$. Therefore, given that $Y_{1}\perp\!\!\!\perp Y_{0}$ we have that $(Y_{1}-Y_{0})\sim\mathcal{N}(\mu_{1}-\mu_{0},\sigma_{1}^{2}+\sigma_{0}^{2}).$ If larger values are ’better’, then the win probability of $Y_{1}$ vs. $Y_{0}$ (see for example Gasparyan et al.[24]) is
@@
as the theoretical version of Cohen’s d.[25] Hence, the win probability is a standardized effect measure as opposed to non-standardized effect measures such as difference in means or medians. Consequently, win statistics are rather a measure of discrimination between two distributions. In the case of continuous distributions, the probability of a tie is 0, and thus the WR and WO are the same and are defined as
@@
The WR’s dependency on variation, particularly when measurement error is a substantial contributor to the variation in a continuous variable, closely parallels the scenario encountered when dichotomizing continuous endpoints for ’responder analysis.’. This is thoroughly examined in Abugov et al.[26], where they conclude that "In other words, large responder effect benefits do not necessarily imply large clinical benefits and small responder effect risks do not necessarily imply minimal clinical risk. A large responder benefit may reflect an unimportant clinical benefit amplified by a small standard deviation, and a small responder safety risk may reflect a clinically important change minimized either by a large standard deviation… ". Similarly, the draft FDA guidance on obesity and overweight[27] reflects this concern, stating that "…responder analyses for the evaluation of weight reduction are generally not recommended.".
Note that this is a property of the theoretical win statistics, that is, the estimand itself. In what follows, we will explore the impact of different sources of variability on the WR, including impact of the variability of the measurement error.
@@ 3.2 Slope example
Figure 2 shows how the WR depends on the population slope variability for each of the indicated analysis methods as well as for the latent true slopes for a two-year trial in which GFR is measured every 3 months, the between-patient slope SD ($\sigma_{s}$) is 3 ml/min/1.73m2/yr, the treatment attenuates the mean slope from -3 ml/min/1.73m2/yr in the control group to -2 ml/min/1.73m2/yr in the treatment group, and the within patient error standard deviation ($\sigma_{e}$) is 5.18 ml/min/1.73m2. The within patient error standard deviation is typical of past RCTs with mean baseline GFR of approximately 40 ml/min/1.73m2, see eg Greene et al.[28] The LSME and MC methods underestimates the WR in this case by 9% and 13% respectively. The WR decreases as a function of the between-patient slope standard deviation for both the true slopes and for each of the methods of slope estimation. For any given between-patient slope standard deviation, we see the expected result; both methods underestimates the compared to the ’true’ slopes, and the LSME performs better than the MC method.
Figure 2: WR vs between-subject slope standard deviation for the LSME and MC methods, and the true value
@@
Figure 3: WR vs Follow-up time for the LSME and MC methods, and the true value
Figure 4: WR vs number of GFR measurements for the LSME and MC methods, and the true value
@@ 3.3 Non-collapsibility
Daniel et al. (2020) discussed the concept of non-collapsibility of effect measures.[29] In short, a treatment effect measure is said to be non-collapsible if the inclusion of a covariate in the model associated with the outcome will change the target of the treatment effect to be estimated. A non-collapsible effect measure, such as the odds ratio or the hazard ratio, will have different conditional and marginal estimates. A WR or a WO, in general, is also a non-collapsible effect measure. This can be illustrated by the following example for a WR based on a continuous variable, e.g. GFR.
@@
Figure 5: Normal density plots for two treatments across two strata, illustrating the distribution patterns.
The example is similar to the one given for the Odds ratio in the FDA guidance on adjusting for covariates in randomized clinical trials.[30]
In this example the population consists of two equally sized strata. The endpoint is normally distributed with population means given in the table. The standard deviation for both strata and treatments are given as $\sigma_{sj}=\sigma=5$. The strata specific WRs calculated by means of equation (5) are 3.17 for both strata. The marginal WR for the entire population calculated by means of equation (6) is 2.18. The distributions are illustrated in  5, here it can be seen that when doing comparisons across strata the resulting WR will be "diluted" from the cross-strata comparisons. Note that the distribution in the ‘combined’ population is bimodal. A stratified WR would yield the same point estimate as the strata-specific WRs, since any weighted average results in the same value. As for the odds ratio and hazard ratio, the non-collapsibility of the WR makes it problematic as a causal effect measure and complicates its transportability between different populations. This issue is particularly pronounced in meta-analyses, where combining results from different studies becomes challenging. More details on causality and non-collapsibility can be found in Didelez & Stensrud (2021).[31]
@@ 4 Discussion
In this study, we have explored some of the theoretical and practical implications of using the Win Ratio (WR) in the context of Hierarchical Composite Endpoints (HCEs) in clinical trials, particularly focusing on chronic kidney disease (CKD). Our findings highlight several critical issues that warrant careful consideration, including aspects that have received little attention in the literature.
@@
The use of WR in HCEs raises several theoretical concerns. One of the primary issues is the challenge of defining a suitable estimand that accurately reflects the clinical question being addressed. The WR, while useful in certain contexts, does not always provide a clear interpretable measure of the treatment effect. However, in some trials, it may be infeasible to address the most relevant clinical questions, such as overall long-term survival, necessitating compromises that hinder rigorous adherence to the principles of the estimand.
The actual value of the WR depends on the variance of any continuous components included in the HCE. This dependence implies that the WR primarily serves as a measure of discriminatory character between distributions of active and control groups rather than a pure effect measure. The example of GFR slope illustrates that these issues are substantial in realistic setups - we observed a bias of over 10% in realistic settings. The variance of the continuous components can markedly influence the WR.
The WR is a non-collapsible measure. This property complicates any causal interpretation of the WR. The issue is particularly problematic when comparing results across different trials or conducting meta-analyses.
@@
The WR is a population-level effect measure, which compares the outcomes of different subjects across treatment groups. This can lead to different conclusions than the more natural, but non-identifiable, individual-level effect measures, which compare potential outcomes within the same subject. Hand’s paradox illustrates this discrepancy between these two. The toy example presented in table 1 is intentionally crafted to highlight the distinction between these. It is extreme in that one patient experiences a significant detrimental effect from the active treatment, while the two other patients experience only a minor beneficial effect. In this scenario one could question the relevance of any overall effect estimate. The main point of this example is to demonstrate that the individual-level measure provides a more relevant causal measure; however, this is not identifiable, and the two measures are not guaranteed to be even similar. Recently, Even and Josse (2025)[32] suggested that a matching approach based on a nearest neighbour algorithm, which can be seen as an extreme form of stratification, could be a way to approximate the non-identifiable individual-level effect measure. However, Pocock et al. (2024)[33] do not recommend the stratified WR in general
While the use of WR in clinical trials should not be discouraged in establishing the treatment effect, it is important to acknowledge the difficulties in interpreting the treatment effect. We propose considering WR as a discriminatory measure between distributions in active and control groups, similar to non-parametric tests. One such application is being presented in Hartley et al (2025)[34]. Although WR offers a novel approach to analyzing HCEs, it presents challenges. Researchers must consider its dependence on variance and non-collapsibility, which complicates the definition of a relevant estimand and makes causal interpretation difficult.

### v3

@@ 4 Variance dependence
In this section, we consider the impact of the variance of a single normally distributed continuous random variable on the win statistics when there is only one continuous component present in the HCE. This simplification helps to clearly illustrate the concept, although the effects will persist even when more components are included, albeit potentially diluted.
@@ 4.1 The impact of the overall variance
Let $Y_{0}\sim\mathcal{N}(\mu_{0},\sigma_{0}^{2})$ and $Y_{1}\sim\mathcal{N}(\mu_{1},\sigma_{1}^{2})$ be two independent normally distributed random variables, where $Y_{i}$ denotes the response under treatment $i=0,1$. Therefore, given that $Y_{1}\perp\!\!\!\perp Y_{0}$ we have that $(Y_{1}-Y_{0})\sim\mathcal{N}(\mu_{1}-\mu_{0},\sigma_{1}^{2}+\sigma_{0}^{2}).$ If larger values are ’better’, then the win probability of $Y_{1}$ vs. $Y_{0}$ (see for example [16]) is
@@
as the theoretical version of Cohen’s d [32]. Hence, in this case, the win probability is a standardized effect measure as opposed to non-standardized effect measures such as difference in means or medians. Consequently, win statistics are rather a measure of discrimination between two distributions. In the case of continuous distributions, the probability of a tie is 0, and thus the WR and WO are the same and are defined as
@@
The WR’s dependency on variation, particularly when measurement error is a substantial contributor to the variation in a continuous variable, closely parallels the scenario encountered when dichotomizing continuous endpoints for ’responder analysis.’. This is thoroughly examined in [33], where they conclude that "In other words, large responder effect benefits do not necessarily imply large clinical benefits and small responder effect risks do not necessarily imply minimal clinical risk. A large responder benefit may reflect an unimportant clinical benefit amplified by a small standard deviation, and a small responder safety risk may reflect a clinically important change minimized either by a large standard deviation… ". Similarly, the draft FDA guidance on obesity and overweight [34] reflects this concern, stating that "…responder analyses for the evaluation of weight reduction are generally not recommended.".
Note that this is a property of the theoretical win statistics, that is, the estimand itself. In what follows, we will explore the impact of different sources of variability on the WR, including impact of the variability of the measurement error. A similar exposition can be found in [35].
@@ 4.2 Slope example
Figure 2 shows how the WR depends on the population slope variability for each of the indicated analysis methods as well as for the latent true slopes for a two-year trial in which GFR is measured every 3 months, the between-patient slope SD ($\sigma_{s}$) is 3 ml/min/1.73m2/yr, the treatment attenuates the mean slope from -3 ml/min/1.73m2/yr in the control group to -2 ml/min/1.73m2/yr in the treatment group, and the within patient error standard deviation ($\sigma_{e}$) is 5.18 ml/min/1.73m2. The within patient error standard deviation is typical of past RCTs with mean baseline GFR of approximately 40 ml/min/1.73m2, see eg [36]. The LSME and MC methods underestimates the WR in this case by 9% and 13% respectively. The WR decreases as a function of the between-patient slope standard deviation for both the true slopes and for each of the methods of slope estimation. For any given between-patient slope standard deviation, we see the expected result; both methods underestimates the compared to the ’true’ slopes, and the LSME performs better than the MC method.
Figure 2: WR vs between-subject slope standard deviation for the LSME and MC methods, and the true value, illustrating both the bias of the estimates, and the true/estimated WR’s dependence on $\sigma_{s}$
@@
Figure 3: WR vs Follow-up time for the LSME and MC methods, and the true value, illustrating both the bias of the estimates, and the true/estimated WR’s dependence on follow-up time.
Figure 4: WR vs number of GFR measurements for the LSME and MC methods, and the true value, illustrating the bias of the estimates, and the estimated WR’s dependence on the number of measurements.
@@ 4.3 Non-collapsibility
Daniel et al. (2020) discussed the concept of non-collapsibility of effect measures [37]. In short, a treatment effect measure is said to be non-collapsible if the inclusion of a covariate in the model associated with the outcome will change the target of the treatment effect to be estimated. A non-collapsible effect measure, such as the odds ratio or the hazard ratio, will have different conditional and marginal estimates. The WR, WO, and NB are in general also non-collapsible effect measures, see eg. [38]. To demonstrate this concept, we provide a simple example of the win ratio that utilizes a normally distributed variable, e.g. GFR, as its sole component.
@@
Figure 5: Normal density plots showing the distribution patterns for two treatments across two strata. Intra-strata comparisons indicate that the new drug typically shows a consistent winning trend in both strata; however, this consistency does not hold true for cross-strata comparisons
The example is similar to the one given for the Odds ratio in the FDA guidance on adjusting for covariates in randomized clinical trials [39].
In this example the population consists of two equally sized strata. The endpoint is normally distributed with population means given in the table. The standard deviation for both strata and treatments are given as $\sigma_{sj}=\sigma=5$. The strata specific WRs calculated by means of equation (5) are 3.17 for both strata. The marginal WR for the entire population calculated by means of equation (6) is 2.18. The distributions are illustrated in  5, here it can be seen that when doing comparisons across strata the resulting WR will be "diluted" from the cross-strata comparisons. Note that the distribution in the ‘combined’ population is bimodal. A stratified WR would yield the same point estimate as the strata-specific WRs, since any weighted average results in the same value. As for the odds ratio and hazard ratio, the non-collapsibility of the WR makes it problematic as a causal effect measure and complicates its transportability between different populations. This issue is particularly pronounced in meta-analyses, where combining results from different studies becomes challenging. More details on causality and non-collapsibility can be found in [40].
@@ 5 Discussion
In this study, we have explored some of the theoretical and practical implications of using the Win Ratio (WR) in the context of Hierarchical Composite Endpoints (HCEs) in clinical trials, particularly with examples from chronic kidney disease (CKD). While the WR offers a structured way to compare treatments, several challenges and limitations complicate its interpretability, particularly regarding its relationship with the estimand framework, causal interpretation, transitivity, variance dependence, and its non-collapsible nature.
The application of win statistics for HCEs as primary analyses has gained acceptance as indicated by two cases of FDA approvals for cardiac amyloidosis (Tafamidis [8], ATTR-ACT ClinicalTrials.gov number, NCT01994889. and Acoramidis [41], ATTRibute-CM ClinicalTrials.gov number, NCT03860935 ). However, it is worth noting that in the case of Tafamidis, there is no mention of the win statistics, only the p-value of the underlying Finkelstein-Schoenfeld method.
@@
The use of WR in HCEs raises several theoretical concerns. One foundational issue arises from defining an appropriate estimand for WR-based analyses. According to the ICH E9 addendum on estimands, the estimand must address the clinical question being investigated while ensuring clarity and interpretability. The WR, while useful in certain contexts, does not always provide a clear interpretable measure of the treatment effect. However, in some trials, it may be infeasible to address the most relevant clinical questions, such as overall long-term survival, necessitating compromises that hinder rigorous adherence to the principles of the estimand.
@@
The WR is a population-level effect measure, which compares the outcomes of different subjects across treatment groups. This can lead to different conclusions than the more natural, but non-identifiable, individual-level effect measures, which compare potential outcomes within the same subject. Hand’s paradox illustrates this discrepancy between these two. The toy example presented in table 1 is intentionally crafted to highlight the distinction between these. It is extreme in that one patient experiences a significant detrimental effect from the active treatment, while the two other patients experience only a minor beneficial effect. In this scenario one could question the relevance of any overall effect estimate. The main point of this example is to demonstrate that the individual-level measure provides a more relevant causal measure; however, this is not identifiable, and the two measures are not guaranteed to be even similar. Recently, [42] suggested that a matching approach based on a nearest neighbor algorithm, which can be seen as an extreme form of stratification, could be a way to approximate the non-identifiable individual-level effect measure.
Non-transitivity is an additional concern, particularly in multi-arm trials or when comparing treatments indirectly. The WR may indicate that treatment A is better than B, B better than C, and yet C better than A, making it impossible to establish a consistent treatment ranking. This paradox highlights the need for caution when interpreting WR results in comparative effectiveness research.
Another important limitation is the WR’s dependence on the variance of continuous components. As shown in our slope example, the magnitude of the WR can be heavily influenced by measurement error and study design characteristics such as follow-up duration and frequency of measurements. This sensitivity reduces the WR’s utility as a pure effect measure and instead positions it more as a measure of discrimination between treatment groups.
An additional key concern is the WR’s non-collapsibility, a property shared with other widely used metrics such as odds ratios and hazard ratios. The non-collapsibility of the WR complicates causal interpretation and can lead to inconsistencies when comparing results across trials or conducting meta-analyses. This issue becomes particularly critical when the population differs significantly.
Although the WR holds promise in its ability to evaluate complex composite endpoints, the aforementioned limitations underscore the importance of complementary analyses. Key recommendations include: (1) defining fixed follow-up periods; (2) dissecting individual HCE components and employing visualizations to clarify the drivers of observed WR values, see e.g. [43]; and (3) avoiding the combination of components with vastly differing levels of severity.
In summary, while the WR introduces an innovative and versatile framework for summarizing multifaceted outcomes, its limitations must be carefully considered. We propose considering WR as a discriminatory measure between distributions in active and control groups, similar to non-parametric tests. One such application is being presented in [44] who use win statistics to improve commit-to-Phase-3 decision-making in oncology.
6 Acknowledgment
The authors would like to acknowledge the use of generative AI (ChatGPT, OpenAI o3 Mini) for language enhancements in the preparation of this work.